### v2

# Split and Rephrase with Large Language Models

Abstract
The Split and Rephrase task, which consists in splitting complex sentences into a sequence of shorter grammatical sentences, while preserving the original meaning, can facilitate the processing of complex texts for humans and machines alike. In this work, we describe an approach based on large language models, which improves over the state of the art by large margins on all the major metrics for the task, on publicly available datasets. We also describe results from two human evaluations that further establish the significant improvements obtained with large language models and the viability of the approach. We evaluate different strategies, including fine-tuning pretrained language models of varying parameter size, and applying both zero-shot and few-shot in-context learning on instruction-tuned language models. Although the latter were markedly outperformed by fine-tuned models, they still achieved promising results overall. Our results thus demonstrate the strong potential of different variants of large language models for the Split and Rephrase task, using relatively small amounts of training samples and model parameters overall.
Split and Rephrase with Large Language Models
[…]
**footnotetext: Equal contribution.
1 Introduction
Transforming complex sentences into a sequence of shorter sentences, while preserving the original meaning, is an important task which can facilitate the processing of complex texts for humans and machines alike. It is, for instance, an important operation in text simplification Zhu et al. (2010); Narayan and Gardent (2014); Shardlow (2014), where simpler split sentences can provide content that is easier to read and comprehend.
This task, commonly referred to as Split and Rephrase (SPRP) (Narayan et al., 2017), is typically performed via dedicated transformation models, e.g. sequence-to-sequence modelling with a copy mechanism (Aharoni and Goldberg, 2018) or graph-based neural segmentation (Gao et al., 2021). The main limitation of this type of approach is the dependency on training datasets of paired complex and simple sentences, which are scarce across languages and domains. As an alternative, Kim et al. (2021) exploit one-to-many alignments in parallel corpora and pivot machine translation (MT) to create an aligned corpus of complex and split sentences, which can be exploited with sequence-to-sequence models. For this type of approach, the main bottleneck is the dependency on parallel corpora with sufficient many-to-one alignments, which can also be scarce overall.
Pretrained large language models (LLMs), such as the GPT models (Radford et al., 2018) based on the Transformer architecture (Vaswani et al., 2017), have demonstrated their impressive potential to generate fluent text from natural language prompts, opening new research paths for a large number of downstream tasks. Their potential has been studied, for instance, for MT Bawden and Yvon (2023); Zhang et al. (2023), for translation quality estimation Kocmi and Federmann (2023), or for the generation of synthetic sentence similarity datasets via hand-crafted prompts Schick and Schütze (2021). We follow the spirit of these approaches and evaluate the potential of LLMs for the controlled generation of sequences of meaning-preserving short sentences from complex ones.
We measure the ability of different pretrained LLMs to perform the SPRP task, using task-specific prompts that take a complex sentence as input and probe the language model for a continuation in the form of split sentences. We evaluate our LLM-based approach on publicly available English datasets and explore controlled generation under zero shot, few shot in-context learning, and fine-tuning scenarios. Our analysis includes the use of LLMs with varying number of parameters, to establish minimal and optimal configurations for the SPRP task. We also evaluate the accuracy of LLMs trained only on the next token generation task and models further tuned for instruction completion. Finally, we measure the impact of training data size on SPRP accuracy of fine-tuned LLMs.
Overall, our results demonstrate the strong potential of pretrained LLMs for SPRP, outperforming the state of the art by large margins across multiple metrics. We also report results of two manual evaluations, a comparative 3-way ranking task and a qualitative evaluation of LLM-generated SPRP output, both of which further establish the gains and output quality achievable by our approach.
2 Related Work
The Split and Rephrase task was first proposed by Narayan et al. (2017), who created the WebSplit dataset from RDF subject-predicate-object semantic triples and also evaluated initial benchmarks using different types of models, including statistical MT models and encoder-decoder LSTMs with attention Bahdanau et al. (2015). Aharoni and Goldberg (2018) reported that the initial baselines suffered from data memorisation and lack of RDF relation types in the training set. They proposed a new data split and a new benchmark using LSTMs using attention and a copy mechanism (Gu et al., 2016), to bias the model towards copying the source. Botha et al. (2018) argued that WebSplit was limited by its small vocabulary and unnatural linguistic expressions, and presented WikiSplit, a corpus created using Wikipedia data that takes advantages of edits made on a page to split or fuse sentences. They showed that models trained with WikiSplit dataset obtained better BLEU Papineni et al. (2002) scores than those trained on WebSplit. However, they also noted that WikiSplit is a relatively noisy dataset, not suited for evaluation.
In Das et al. (2018), simple sentences are generated from complex and compound sentences by leveraging dependency parsing information, achieving promising results. Niklaus et al. (2019) employed handcrafted rules, recursively applied to complex sentences to generate simpler ones, which they reported to be beneficial in downstream tasks. Recent approaches have exploited neural modelling, and in particular the Transformer architecture (Vaswani et al., 2017). Thus, Malmi et al. (2019) proposed a tagging approach using a BERT model (Devlin et al., 2019) as encoder and a single Transformer decoder layer. Stahlberg and Kumar (2020) also propose a tagging model, at the span level instead of the token level, obtaining improvements with a Pointer Network (Vinyals et al., 2015) for end of span prediction and a two-layer Transformer decoder. Berlanga Neto and Ruiz (2021) introduced a cross-lingual model based on part-of-speech tags, GRU sequence-to-sequence modelling, and BERT for next word prediction. Recently, AlAjlouni and Li (2023) reported improvements via transfer knowledge of data generated by the DISSIM rule-based system (Niklaus et al., 2019) to fine-tune an encoder-decoder T5 model .
In this work, we compare our LLM-based approach with two recent approaches and associated corpora. First, we address the work of Kim et al. (2021), who proposed BiSECT, a method and corpus based on extracting 1-2 alignments from parallel bilingual corpora and machine translating the sentences into the same language. They showed that BiSECT was of higher quality than WikiSplit, and that sequence-to-sequence models trained on it achieved better results on the SPRP task overall. Additionally, we compare our work to Gao et al. (2021), who proposed the ABCD framework where SPRP is performed via neural graph segmentation, according to the number of predicates in the complex sentence, showing that this approach approached or outperformed the performance of parsing-based and encoder-decoder models. They also provided two new datasets for the SPRP task: MinWiki, a reduced version of WikiSplit where only pairs of complex and simple sentences with a matching number of predicates are kept, and DeSSE, a corpus of students’ opinion essays in social science consisting of complex sentences and corresponding splits; we used the latter in our experiments, in addition to BiSECT.
Finally, we used the Wiki-BM corpus described in Zhang et al. (2020a), who found that the standard WebSplit dataset contained easily identifiable syntactic cues, which even a simple rule-based approach could exploit to match the performance of state-of-the-art models. Their benchmarks feature higher syntactic diversity to overcome this issue and are thus suited for our qualitative evaluation of LLM-based SPRP hypotheses.
3 Methodology
Our approach is based on a simple idea: prompt a LLM to generate sequences of sentences from a single complex sentence. We explored two variants to control text generation for these tasks.
We first fine-tuned pretrained LLMs on task-specific prompts, querying for complex sentence split. The objective in this scenario is twofold: (i) measure the performance of LLMs that have not been previously fine-tuned for instruction-based task completion, and (ii) evaluate the accuracy of LLMs depending on their number of parameters. The latter point is relevant in particular to evaluate minimal and sufficient needs in terms of computational resources for the task.
Additionally, we evaluate scenarios where we directly prompt LLMs that have been previously tuned on instruction datasets. The goal in this case is to measure the ability of this type of readily trained models to perform the two tasks at hand, with or without positive examples. Under this scenario, we explore both zero-shot and in-context few-shot task completion, by directly prompting the model for completion in the first case, and providing a limited number of examples within the prompt in the latter case.
4 Experimental Setup
4.1 Corpora
To train and evaluate our models we selected two recent datasets in English for the SPRP task: DeSSE (Gao et al., 2021) and BiSECT (Kim et al., 2021). [^1] We use the datasets made available in the following repositories, as of April 2023: https://github.com/serenayj/DeSSE and https://github.com/mounicam/BiSECT. Statistics for each corpus are provided in Table 1, in terms of number of aligned pairs of complex and split sentences. We indicate the figures for the original dataset (Original) and the versions we employed in this work (Selected).
Corpus Version Train Dev Test
DeSSE Original 13,199 42 790
DeSSE Filtered 6,432 2,000 504
BiSECT Original 928,440 9,079 583
BiSECT Filtered 6,432 2,000 580
Table 1: DeSSE and BiSECT English corpora statistics
For testing, we filtered all pairs of complex and split sentences in each corpus where the latter consisted of only one sentence, instead of multiples ones. There were 286 such instances in the DeSSE corpus, and 3 in BiSECT. We used the filtered version of the test datasets for all experiments reported below, as we aimed to evaluate model performance strictly on pairs of complex sentences and corresponding sequences of multiple sentences.
In terms of training data, we also filtered the DeSSE training set along the same lines, and used the resulting dataset for fine-tuning. For BiSECT, we randomly sampled the same amount of training data, for comparison purposes. As development data, since the original sets for DeSSE were rather small (42 examples), we extracted 2,000 pairs from the training set for both datasets.
The two corpora differ in several respects. First, in terms of size, BiSECT is orders of magnitude larger in terms of training and development datasets, although the filtered version of both test sets are of comparable size. In terms of domain, DeSSE contains students opinion essays collected from a social science class and is therefore less varied in terms of topics than BiSECT, which was created from a large variety of parallel corpora. Finally, the average length of the complex sentences, computed in terms of tokens over tokenised sentences, is also markedly different: for the training datasets, it amounts to an average of 40.14 tokens in BiSECT, contrasting with 25.25 in DeSSE. These differences are maintained for the test sets, with average lengths of 39.55 for BiSECT and 24.10 for DeSSE.
4.2 Models
Pretrained language models.
As pretrained LLMs, we used the Pythia suite (Biderman et al., 2023), a collection of Transformer-based decoder-only autoregressive generative language models available in different parameter size and trained on the Pile dataset (Gao et al., 2020). We selected the following variants for our experiments, in terms of number of model parameters: 410M, 1B, 1.4B, 2.8B, 6.9B and 12B. We fine-tuned each of these models with LoRA (Hu et al., 2022), using default parameters and the following prompt, where $\alpha$ is a complex sentence and $\beta$ its corresponding split sentences, sampled from the selected portions of the DeSSE and BiSECT training sets: [^2] See Appendix A for more details on LLM training.
  • INSTRUCTION: Split the sentence below into separate sentences.
  • INPUT: $\alpha$
  • RESPONSE: $\beta$
Although other types of prompts are conceivable for the task, with different instructions or markers to denote the components of a prompt, preliminary results with prompt variants did not lead to significantly different results and we left further exploration on this matter for future research. At inference time, the response is empty and to be completed by the model.
Instruction-tuned language models.
As LLMs readily tuned for instruction-based prompting, we selected Dolly 2.0, [^3] https://huggingface.co/databricks/dolly-v2-7b with 7B parameters, which is based on the 6.9B Pythia language model and was tuned on 15K instructions. [^4] https://huggingface.co/datasets/databricks/databricks-dolly-15k This model allowed for a direct comparison between the fine-tuned LLM and the instruction-based derived model. Additionally, we selected the Alpaca-Lora model with 7B parameters, based on LLaMA Touvron et al. (2023) and fined-tuned on the Alpaca instruction dataset, [^5] https://github.com/gururise/AlpacaDataCleaned representative of a model derived from a different LLM and set of instructions.
As indicated in Section 3, we used two different approaches to probe these model variants. For zero-shot probing, we used the prompt patterns described above, leaving the response empty for the model to complete. For in-context learning, we used a few-shot approach based on $n$ examples prepended to the prompt. Thus, each query is composed of $n$ concatenated prompts that follow the previously described pattern, along with the last prompt where the response is left empty. For our experiments, we selected the first 5 examples from the development set as examples, although other variants are of course possible.
Figure 1: Split and Rephrase results on the DeSSE (left) and BiSECT (right) test sets, with Pythia model variants fine-tuned over data from the DeSSE (-DE) or BiSECT (-BI) training data.
4.3 Evaluation
Following Kim et al. (2021), we used three separate metrics to evaluate our models. SARI (Xu et al., 2016), the standard simplification metric, and more specifically the version computed with the EASSE toolkit (Alva-Manchego et al., 2019) [^6] https://github.com/feralvam/easse, which provides the proportion of ngrams that are added (add in what follows), kept (keep) or deleted (del). As a second metric, we used the normalised BertScore (hereafter, BESC) (Zhang et al., 2020b), [^7] https://github.com/Tiiiger/bert_score which has proved successful to measure meaning preservation. BLEU scores, computed with sacreBLEU (Post, 2018), are reported as well, as a precision measure of n-gram overlap between hypotheses and references.
We also included an additional metric for the task to indicate output compliance (CPL), i.e. the percentage of output sequences that include at least two separate sentences, to directly identify unsplit output. This metric was computed with a simple approximation, counting the number of end of sentence punctuation marks in the output sentence, including period, question and exclamation.
We compare our results with those obtained using the BiSECT and ABCD approaches of Kim et al. (2021) and Gao et al. (2021), respectively, which represent the state of the art for the selected corpora. For the BiSECT approach, we used the publicly shared output available for their model. [^8] https://github.com/mounicam/BiSECT/tree/main/outputs/
bisect/our_model_bisect.txt
For the ABCD graph segmentation approach, we recomputed the output of their models using the publicly available code. [^9] https://github.com/serenayj/DeSSE
Figure 2: Split and Rephrase results on the DeSSE (left) and BiSECT (right) test sets, for instruction-tuned Alpaca and Dolly model variants. Models were applied either in a zero-shot fashion (-ZS), or in a few-shot manner (-FS) with 5 examples sampled from the DeSSE (-DE) or BiSECT (-BI) datasets.
5 SPRP with Large Language Models
5.1 Fine-tuned LLM Variants
We first evaluated the Pythia LLM variants fine-tuned on the selected DeSSE and BiSECT training data. The results are shown in Figure 1.
In terms of parameter size, on all three metrics there is a slight tendency towards increased accuracy as the number of parameters increases, although the differences are relatively minor overall. The sole exception are the BLEU results with models fine-tuned on BiSECT and evaluated on DeSSE. In this specific scenario, the behaviour was also more erratic overall on the SARI metric, with marked drops for the 1B model.
Fine-tuning on DeSSE data led to the highest scores on all three metrics, with an average of 84.96 for BESC, 72.63 for BLEU and 56.98 for SARI, against average scores of 74.45, 38.88 and 33.27, respectively, for models fine-tuned on BiSECT data. As previously mentioned, the DeSSE training data feature both shorter sentences and lesser domain variation, which may lead to an easier modelling of the patterns and operations to be replicated at inference time. We leave a more precise analysis of these factors for future work, as it is currently relatively unclear how to precisely determine which characteristics of the fine-tuned samples may impact controlled text generation on the selected tasks. From these results, a preliminary practical conclusion might be to favour simpler and shorter data samples to fine-tune LLMs for the SPRP task.
Domain shift was detrimental in both cases. For models tested on DeSSE, the drop from in-domain to out-of-domain fine-tuning averaged to 6.82 points on BESC, 16.53 on BLEU, and 20.67 on SARI. Although less marked, models tested on BiSECT while fine-tuned on DeSSE experienced an average drop of 2.39, 3.02 and 5.55, on BESC, BLEU and SARI, respectively. The relatively larger decrease in scores with DeSSE-tuned models may be attributed to the application of these models on BiSECT test sets, which contain longer sentences on average. However, this drop still resulted in markedly higher scores compared to the BiSECT-tuned variants on the same dataset.
5.2 Instruction-tuned Models
The zero-shot and in-context few-shot results with the Alpaca and Dolly models on the SPRP task are shown in Figure 2. The first notable result is the overall higher scores obtained with Alpaca models, on all metrics, when compared to equivalent Dolly variants. The differences were not marginal as they amounted to average differences, between comparable model sizes, of 13.06, 18.57 and 13.93 points in terms of BertScore, BLEU and SARI, respectively, on the DeSSE test set. On the BiSECT test, the average differences between comparable models amounted to 8.09, 7.23 and -0.97, with only the SARI metric leading to comparable results on this test set. The choice of a specific pretrained instruction-tuned LLM for the SPRP task thus seems of particular importance, despite their relative similarity in terms of model architecture, training objectives and dataset variety.
Few-shot variants outperformed their zero-shot alternatives across the board, although it is worth noting that the Alpaca zero-shot model outperformed the Dolly few-shot model overall. As was the case for the fine-tuned Pythia variants, the results on the BiSECT test set were also significantly lower overall than on the DeSSE test, due in part to the higher complexity of this specific test set, which features significantly longer sentences on both the complex and split sides.
5.3 Impact of Training Data Size
The results described in the previous sections for the best performing models were obtained with approximately 6,000 training data points. In this section, we aim to evaluate LLM accuracy on the SPRP task with either larger volumes of training data, testing the ceiling of the approach, or smaller amounts to estimate minimal volumes of training data for this method. We used the fine-tuned Pythia-6.9B models for these experiments, per the previously discussed rationale. On the BiSECT corpus, which features larger amounts of data, we performed our experiments in both directions, increasing or decreasing the size of the training datasets. For the smaller DeSSE corpus, since we had already used the complete dataset, we only explored smaller training data partitions.
The results are shown in Table 2.101010In all tables, best results are shown in bold. In both cases, scores increased across metrics as training size increased, although by smaller increments from the 3K mark upwards. Best results are split between the two largest training datasets on both BiSECT and DeSSE, indicative of a potential ceiling in terms of improvements from training data augmentation. It might also be the case that the observed plateauing resulted from a lack of variety in the added training data, but verifying this hypothesis was beyond the scope of this work.
The two domains displayed noteworthy differences. Although using all the available training data resulted in better SARI and compliance scores on DeSSE, the scores obtained with 3K were closer to the best results than to those obtained with 2K training data points. For BiSECT on the other hand, compliance is under 67% for models trained with 4K data points or less, with a significant increase to 82% when using 6K and a stabilisation above 90% with training datasets of 14K or more samples. These results are in line with the previously discussed complexity of the BiSECT datasets, which feature long sentences, machine-translated content and a variety of domains which makes the SPRP task more difficult to model, compared to the relatively simpler DeSSE datasets.
Train Size BESC BLEU SARI CPL
BI 1k 70.69 35.90 21.78 10.34
BI 2k 74.31 36.80 21.23 5.17
BI 3k 75.61 40.00 32.43 52.93
BI 4k 75.82 40.70 35.69 66.03
BI 6k 76.12 41.30 38.93 81.55
BI 14k 76.04 41.60 41.93 90.69
BI 30k 76.32 42.40 45.06 95.17
BI 62k 76.74 42.60 43.84 93.79
DE 1k 69.21 48.60 32.73 13.89
DE 2k 82.26 73.90 59.99 73.21
DE 3k 88.25 81.10 67.33 85.52
DE 4k 88.80 81.60 68.58 89.09
DE 6k 88.78 81.80 70.39 94.64
Table 2: Impact of training data size with the Pythia-6.9B model on BiSECT (BI) and Desse (DE)
6 Comparative Evaluations
Method Test BESC BLEU SARI add keep del CPL
pythia-6.9B-DE DeSSE 88.78 81.80 69.47 45.75 92.26 70.39 94.64
alpaca-7B-ZS DeSSE 75.46 64.60 56.44 23.14 89.35 56.82 83.93
alpaca-7B-FS-DE DeSSE 84.55 74.90 56.08 27.76 88.92 51.56 68.65
ABCD-bilinear DeSSE 66.27 57.20 52.54 17.63 81.04 58.96 98.41
ABCD-mlp DeSSE 78.68 67.40 57.67 22.04 87.90 63.06 89.09
pythia-6.9B-BI BiSECT 76.12 41.30 38.93 12.92 63.59 40.27 81.55
alpaca-7B-ZS BiSECT 69.32 36.20 29.82 5.08 60.39 23.99 88.62
alpaca-7B-FS-BI BiSECT 73.17 36.80 35.82 8.57 60.46 38.43 65.86
BiSECT seq2seq BiSECT 72.51 36.80 33.21 7.34 59.91 32.40 99.48
Table 3: Comparative metrics results for the Split and Rephrase task on the DeSSE and BiSECT test sets.
6.1 Metrics Results
From the results in the previous sections, we selected a subset of models for comparison with the ABCD and BiSECT approaches, namely: the 6.9B Pythia models, which obtained only slightly lower results on all metrics than the 12B variant and features a comparable amount of parameters to the instruction-tuned LLMs; the Alpaca zero-shot model, representative of scenarios where no data are available for fine-tuning or in-learning; and the Alpaca few-shot model using either BiSECT and DeSSE examples. For the ABCD approach, we used both the bilinear and MLP variants described in Gao et al. (2021). The comparative results are shown in Table 3.
On both test sets, the fine-tuned Pythia models with 6.9B parameters obtained significantly better results than all alternatives. On DeSSE, the pythia-6.9B-DE model outperformed the best ABCD model (ABCD-mlp) by 10.1, 14.4, 11.8 and 5.55 points in terms of BESC, BLEU, SARI and CPL, respectively; the ABCD-bilinear model obtained the best compliance score but the lowest results on the other metrics. On the BiSECT test sets, the pythia-6.9B-BI model outperformed the BiSECT method by 3.61, 4.5 and 5.72 points on BESC, BLEU, SARI, respectively; the BiSECT model only obtained better results in terms of CPL. The LLM approach can thus provide more accurate results across the board when a split is generated, but can still lag behind alternatives in terms of enforcing SPRP hypotheses.
Although fine-tuning resulted in the best results overall, it is worth noting that instruction-tuned models prompted with only 5 in-context examples obtained competitive results across the board. Their main drawback involved the lower compliance scores, compared to all other alternatives. The fine-tuned Pythia models were thus shown to be a better alternative when training datasets of sufficient size are available, but the instruction-based models could provide a reasonable alternative. The zero-shot results were also competitive against the ABCD and BiSECT approaches, although they obtained the lowest scores overall. In the few-shot approach, one limiting factor is the maximum context size, whose exploration we left for future research.
6.2 Manual Ranking Results
Metrics such as BLEU have important limitations for the SPRP task, displaying low correlation with human judgements in some experiments Zhang et al. (2020a). Although the results of the previous section show large improvements across multiple metrics, indicative of marked improvements achieved by LLM-based SPRP, we conducted an additional manual comparative evaluation. For this specific evaluation, human evaluators were presented a source sentence and two SPRP hypotheses, which they had to rank in a 3-way evaluation, with either one of the hypotheses considered better, or both considered equally correct or incorrect.
For this task, we adapted the Appraise environment Federmann (2018), typically used for machine translation evaluation campaigns and asked 11 evaluators to perform two sub-tasks, each involving 100 samples randomly selected: compare model hypotheses from ABCD-mlp and Pythia-6.9B fine-tuned, on the DeSSE test set, and compare hypotheses from Pythia-6.9B fine-tuned and the BiSECT model on the BiSECT test set. Evaluators were asked to compare the hypotheses according to SPRP criteria of compliance, sensicality, grammatically, missed facts and omitted facts. They were also asked to ignore minor capitalisation and tokenisation issues, which were typically present in the output of the ABCD and BiSECT models, as detected in preliminary data inspection. [^11] See Appendix D for details on the experimental protocol. The results are shown in Table 4.
S P > S P = S P < S $\alpha$
ABCD 70.27% 28.55% 1.18% 0.52
BiSECT 43.09% 38.45% 18.45% 0.49
Table 4: 3-way ranking results constrasting hypotheses from Pythia-6.9B (P) and ABCD or BiSECT on the DeSSE and BiSECT test sets, respectively, with Krippendorf’s $\alpha$ inter-annotator agreement.
Compared to ABCD, LLM-based hypotheses were largely preferred by human evaluators, at 70.27%, with 28.55% considered equal. On the more complex BiSECT test data, which mainly featured long sentences in the pharmaceutical and medical domain, the results were more balanced, with 38.45% of hypotheses considered equally good or bad. However, LLM hypotheses were still markedly preferred for the remaining hypotheses, at 43.09% against 18.45% of hypotheses selected from the BiSECT sequence-to-sequence model. In both evaluations, inter-annotator agreement was moderate overall, at 0.52 and 0.49, slightly lower on the more complex BiSECT data.
The results obtained in this comparative human evaluation are thus in line with those obtained on automated metrics overall, with marked improvements achieved with LLM-based SPRP.
7 Qualitative Evaluation
To gain further insights on the SPRP capabilities of LLMs, we performed an additional human evaluation, this time centred on assessing the validity of individual SPRP hypotheses on a set of predefined criteria. For this qualitative evaluation, we sampled 100 source sentences from Wiki-BM Zhang et al. (2020a), a manually crafted benchmark created from Wikipedia data, and generated SPRP hypotheses with the Pythia-6.9B model fine-tuned on DeSSE data. The evaluation involved 8 human evaluators, who had to answer 6 questions related to the quality of an SPRP hypothesis considering a source sentence. The questions were adapted from Zhang et al. (2020a), with minor modifications: [^12] See Appendix E for details on the experimental protocol.
  1. 1.
    Is it sensical? (scale of 0-2)
  2. 2.
    Is it grammatical? (scale of 0-2)
  3. 3.
    Does it miss any existing facts? (yes/no)
  4. 4.
    Does it introduce new facts? (yes/no)
  5. 5.
    Does it have splits in the wrong place or no split at all? (yes/no)
  6. 6.
    Should some of its sentences be further split? (yes/no)
Overall, the results, shown in Table 5, were largely positive. Most output was judged sensical, grammatical, with almost no introduction of new facts, and correct splits in over 91% of the cases. The last three were judged in almost perfect inter-annotator agreement, with sensicality slightly below but still above 80% agreement among raters.
Although still featuring scores above 80%, the remaining aspects led to comparatively lower results and inter-annotator agreement, with missed facts and not enough splits in 15.87% and 16.75% of the cases, respectively, and inter-rater agreement between 0.55 and 0.60. Missed facts in particular were somewhat more difficult to judge, according to post-mortem feedback, as some rephrasing cases involved implicit rather than explicit logical connection between sentences, which could be rated more or less strictly. Similarly, the need for additional splits can be viewed as more subjective than other categories such as incorrect splits, as evaluators could view some grammatically correct additional splits as cases of over-splitting. [^13] Examples of LLM output from this evaluation are shown in Appendix B; metrics results on the complete Wiki-BM benchmark are indicated in Appendix C.
It is worth noting that these results were obtained with a model fine-tuned on data from a different domain than Wiki-BM, namely DeSSE, showing that knowledge extracted from one dataset can be effectively transferred to other domains.
Question Answer $\alpha$
Sensical 86.12% / 1.84 0.803
Grammatical 92.62% / 1.91 0.907
No missed facts 84.13% 0.597
No new facts 99.5% 0.996
Correct split 91.13% 0.797
Enough split 83.25% 0.548
Table 5: Qualitative evaluation results with Pythia-6.9B ($\alpha$ = Krippendorf’s alpha inter-rater agreement). Answers for the first two questions are the percentage of 2 and the average score. For the other questions, we report the percentage of yes.
8 Conclusions
In this work, we evaluated a novel approach to the Split and Rephrase task, based on large language models. This approach was tested on two SPRP datasets, DeSSE and BiSECT, improving over the state of the art by significant margins on all metrics. The results of a manual 3-way ranking evaluation further confirmed human preference for LLM-based hypotheses, and a qualitative evaluation on Wiki-BM data further demonstrated the quality of the generated SPRP output.
Of the LLM variants we explored, the fine-tuned models obtained the best results overall, although it is worth noting though that the instruction-tuned LLMs outperformed or matched the state of the art using only 5 in-context examples, on all metrics but compliance. They might thus be a reasonable alternative using hand-crafted examples in languages where training corpora are lacking for the task.
Increased data size for fine-tuned models resulted in gradual improvement across metrics, although results stabilised with 6K on BiSECT and 2K on DeSSE. Increasing parameter size resulted in only minor improvements and models with approximately 7B parameters provided the more balanced results overall on the SPRP task.
9 Limitations
The first limitation of our approach is the need for pretrained LLMs for the specific languages in which the SPRP task needs to be performed. Note however that even English-centric LLMs can exhibit multilingual capability Armengol-Estapé et al. (2022), thus mitigating the issue. Additionally, multilingual models such as BLOOM Scao et al. (2022) can provide support for the task; we provide preliminary results in this sense in Appendix F. A more in-depth evaluation of LLM performance on the SPRP task, for languages with limited coverage, would be required to further assess the limitations of the approach.
A second limitation concerns the dependency on training data to fine-tune pretrained LLMs. We established that between 2K and 3K pairs of complex and split sentences might be necessary for the task, with high quality results obtained using between 4K and 6K training pairs, depending on the domain. Although this type of training data can be collected from either parallel multilingual datasets, as in BiSECT, or specific types of essays, as in DeSSE, this still requires dedicated work for new languages. Machine translating existing SPRP corpora with high quality models could provide a basis for fine-tuning in other languages, although a precise evaluation of this approach would be needed (see Appendix F for preliminary results along these lines). We also showed that instruction-tuned pretrained models could provide quality results using only 5 in-context learning pairs, which could be rapidly crafted manually.
10 Ethical Considerations
Our approach involves the use of large language models, whose computational performance is typically higher when deployed in more powerful environments with GPUs, for either training or inference. Under such usage, electric consumption and associated carbon footprint are likely to increase and users of our method under these conditions should be aware of this type of impact. Our approach mitigates these costs via the use of pretrained language models of relatively limited size, and an efficient tuning method such as LoRA Hu et al. (2022) which minimises training time and computational resources.
Appendix A LLM Training and Inference
Fine-tuning was performed with LoRA Hu et al. (2022) using the following default parameters: r=8, alpha=16, dropout=0.05; targeted modules were: Query, Key and Value.
We also used default generation parameters, namely: temperature=0.1; top_p=0.75; top_k=40; num_beans=4; max_new_tokens=128.
Each training sample was provided in the following format, where $\alpha$ is a complex sentence and $\beta$ its corresponding split sentences, sampled from the DeSSE and BiSECT training sets:
  • INSTRUCTION: Split the sentence below into separate sentences.
  • INPUT: $\alpha$
  • RESPONSE: $\beta$
The following text was prepended to each sample prompt: "Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request."
Each training process was performed on 1 A40 GPU with 40GB of RAM. As reference times, fine-tuning took 06:07:59 hours for the Pythia-6.9B model on the BiSECT 6K training dataset and 48 hours over the larger 64K dataset. Fine-tuning the Pythia 12B model took 06:07:59 hours and the smaller 410M model took 00:37:51 hours, on the 6K BiSECT dataset.
Appendix B Split and Rephrase Examples
In Tables 6, 7 and 8, we provide illustrative examples of split and rephrased output obtained with the fine-funed Pythia-6.9B model on Desse, BiSECT and Wiki-BM source sentences, respectively. Sequences from the source text that have been correctly deleted or modified are shown in purple. Sequences that have been correctly added to the output are shown in green. Sequences that have been incorrectly altered in the original sentence are shown in red.
Original Text Split & Rephrased Text
Most people travel to Europe or the Islands and get a taste of the culture in those places.
Most people travel to Europe or the Islands. They get a taste of the culture in those places.
People would’ve been killed, places would have been bombed and there would have been so much going on all at once.
People would’ve been killed. Places would have been bombed. There would have been so much going on all at once.
War is injuring innocent people which is not necessary.
War is injuring innocent people. It is not necessary.
As education continues to reach the poorer areas of our world i truly believe that one day we will all live in TRUE harmony.
Education continues to reach the poorer areas of our world. I truly believe that one day we will all live in TRUE harmony.
This land is still the home to many civilians who rely on the resources in their area to live.
This land is still the home to many civilians. They rely on the resources in their area to live.
Table 6: SPRP examples on the DeSSE dataset
Original Text Split & Rephrased Text
Dialysis does not alter serum doxycycline half-life and thus would not be of benefit in treating cases of overdose.
Dialysis does not change the serum half-life of doxycycline. Thus, it would not be of benefit in the treatment of cases of overdose.
In the first study, type A resistant patients (A-non responders) were randomised to receive placebo or 10, 000 U of NeuroBloc and in the second, type A toxin responsive patients (A-responders) were randomised to receive placebo, 5000 U or 10, 000 U of toxin.
In the first study, type A non-responders were randomized to receive placebo or 10, 000 U of NeuroBloc. In the second study, type A responders were randomized to receive placebo, 5000 U or 10, 000 U of toxin.
Most patients treated with Zevalin experience a severely reduced number of blood cells, which needs some time to normalise again after treatment.
Most patients treated with Zevalin experience a severe reduction in the number of blood cells. It takes some time to return to normal after treatment.
Lisonorm tablet contains 5mg amlodipine/ 10mg lisinopril and is an anti-hypertensive indicated in patients with blood pressure adequately controlled with lisinopril and amlodipine given concurrently at the same dose level.
Lisonorm tablet contains 5 mg of amlodipine and 10 mg of lisinopril. It is indicated for the treatment of hypertension in patients whose blood pressure is adequately controlled with lisinopril and amlodipine.
Table 7: SPRP examples on the BiSECT dataset
Original Text Split & Rephrased Text
She walked out on their wedding , and she currently is single and holds intense anger towards Richard for ruining their wedding, and also became furious to learn she is pregnant, even though Richard told her he got a vasectomy.
She walked out on their wedding. She currently is single. She holds intense anger towards Richard for ruining their wedding. She became furious to learn she is pregnant. Richard told her he got a vasectomy.
Southfort is an unincorporated community in DeSoto County, Florida, United States, located approximately south of Fort Ogden on U.S. Route 17.
Southfort is an unincorporated community in DeSoto County, Florida. It is located approximately south of Fort Ogden on U.S. Route 17.
Parke played for powerhouse youth club FC Delco, and played four years of college soccer at Drexel University, before being drafted by the MetroStars with 60th and last overall pick of the 2004 MLS SuperDraft.
Parke played for powerhouse youth club FC Delco. He played four years of college soccer at Drexel University. He was drafted by the MetroStars with 60th and last overall pick of the 2004 MLS SuperDraft.
Auguste Le Duc started on a bridge between the two islands, which was quickly swept away by the forces of nature.
Auguste Le Duc started on a bridge between the two islands. The bridge was quickly swept away by the forces of nature.
Table 8: SPRP examples on the Wiki-BM dataset
Appendix C Evaluation Metrics on Wiki-BM
In section 7, we used portions of the Wiki-BM corpus (Zhang et al., 2020a) for our qualitative evaluation of LLM-based SPRP output. In Table 9, we provide the results in terms of metrics for the complete Wiki-BM test set, along with results from AlAjlouni and Li (2023), as representative of an alternative method based on fine-tuning an encoder-decoder pretrained T5 model.
The results for our LLM-based approach are in line with those obtained on the DeSSE test set across metrics: slightly lower on BESC, BLEU, and SARI, but slightly higher in terms of compliance (CPL). Our model outperformed the reported scores in AlAjlouni and Li (2023) by 13.4 and 5.1 points on the BLEU and SARI metrics, respectively.
Method BLEU SARI add keep del BESC CPL
pythia-6.9B-DE 72.9 65.1 36.14 91.15 68.10 85.50 98.19
[…]
Table 9: Metrics results on the Wiki-BM benchmark using the Pythia-6.9B model fine-tuned on DeSSE data. Available results from alternative approaches in terms of BLEU and SARI are reported as per the cited work.
Appendix D 3-Way Ranking Evaluation Protocol
Figure 3: Appraise interface adapted for the 3-way ranking SPRP evaluation
The 3-way ranking manual evaluation involved 11 human evaluators, all researchers with at least upper-intermediate knowledge of English, who were not paid for the task. The evaluation was split into two tasks: for one task, the evaluators were asked to rank the output of the fine-tuned Pythia-6.9B model and the output of the ABCD model; for the other task, they had to rank the output of the fine-tuned Pythia-6.9B model and the output of the BiSECT model. Each of the two tasks involved 100 samples.
We adapted the Appraise framework Federmann (2018) for the task, mainly by simplifying the presentation of the options, and the following instructions were provided to complete the task:
  • The task is named Split and Rephrase (SPRP) and consists in splitting a single complex sentence into a sequence of two or more grammatical sentences and rephrasing them as needed to preserve the original meaning without omitting or adding information.
  • You will be shown 100 examples for each task, each example as shown in Figure 3, where:
    • The sentence shown at the top (source) is the original single sentence.
    • The sentences shown as translation A and translation B are SPRP hypothesis which you need to compare.
    • After reading the source sentence and each SPRP hypothesis, you should perform one of the following actions:
        Click on the Better button next to one of the hypothesis if it is better than the other hypothesis according to the criteria below;
        Click on the Equal button if the two hypothesis are both equally correct or both equally incorrect, according to the criteria indicated below.
  • An SPRP hypothesis can be considered correct if:
    • it contains at least two separate sentences (compliance),
    • it contains all the information in the source sentence (no missing facts),
    • it does not contain information that was not present in the source sentence (no new facts),
    • all the sentences in the hypothesis are grammatical (grammaticality),
    • all the sentences in the hypothesis make sense both in isolation and in sequence (sensicality).
  • Ranking criteria:
    • If both hypotheses meet all criteria, select Equal.
    • If neither hypothesis meets any of the criteria, select Equal.
    • If one hypothesis meets all criteria and the other does not, select the former as Better.
    • If one of the hypotheses is markedly better than the other, in the sense that as it meets almost all criteria whereas the other fails to meet most criteria, and only contains minor errors that do not impede comprehension of the original information, you may select the former as Better. In case of doubt, select Equal.
    • When validating an hypothesis, please ignore minor capitalisation errors (e.g., at the beginning of a sentence) and minor tokenisation errors (e.g., punctuation preceded by a space). [^14] This advice was meant to avoid discarding the ABCD or BiSECT hypotheses on these non-SPRP criteria, as they were shown to contain this type of errors in their respective output, upon preliminary data inspection.
System hypotheses are presented in random order by default in the Appraise environmnent.
Appendix E Qualitative Evaluation Protocol
Figure 4: Web interface for the qualitative evaluation
The evaluation involved 8 human evaluators, all researchers with at least upper-intermediate knowledge of English, who were not paid for the task. Each evaluator had to answer 6 questions related to the quality of an SPRP hypothesis considering a source sentence. We randomly sampled 100 sentences from the Wiki-BM corpus Zhang et al. (2020a) and generated SPRP hypotheses with the Pythia-6.9B model fine-tuned on DeSSE data.
The questions were taken from Zhang et al. (2020a) for the most part, with minor modifications: a simplified 0-2 Likert scale instead of their 0-5 scale for the first two questions, as the latter appeared more difficult to use in practice, and a reformulation of question 5 to explicitly include cases where the output was not split.
The samples and associated questions were provided within an ad-hoc Web-based environment developed for the task. The following instructions were provided to the evaluators:
  • Upon entering the Web-based evaluation environment, you will be shown a page as in Figure 4.
  • The topmost sentence is the original source sentence. The sentence below is a Split & Rephrase (SPRP) hypothesis, i.e., a rewrite of the source sentence into a sequence of separate sentences that preserve the meaning of the source sentence.
  • For each pair of source sentence and SPRP hypothesis, you will need to answer all 6 questions in the page. The guidelines for each question are indicated below:
    • Is it sensical? (scale of 0-2)
        Indicate:
          0 if the SPRP hypothesis contains major sensicality errors (the text is barely comprehensible);
          1 if it contains minor sensicality errors (the text is comprehensible but has lost part of the original meaning, e.g., lacking logical connection between sentences);
          2 if there are no sensicality errors.
        Examples:
        • Source: "It was a great evening, until John arrived."
        • Example of 0: "It was an evening. The great John arrived."
        • Example of 1: "It was a great evening. John arrived."
        • Example of 2: "It was a great evening. Until John arrived."
    • Is it grammatical? (scale of 0-2)
        Indicate:
          0 if the SPRP hypothesis contains major grammar errors, which impede text comprehension;
          1 if it contains minor grammatical errors, which barely impede text comprehension;
          2 if there are no grammar errors.
        Examples:
        • Source: "There was a treasure, which was buried on the island."
        • Example of 0: "There was a treasure which. He buried on the island."
        • Example of 1: "There was a treasure. It buried on the island."
        • Example of 2: "There was a treasure. It was buried on the island. "
    • Does it miss any existing fact? (yes/no)
        Indicate:
          Yes if the SPRP hypothesis omits information from the source sentence;
          No if the SPRP hypothesis does not omit information from the source sentence.
        Examples:
        • Source: "There was a treasure on a Greek island, which was never discovered."
        • Example of Yes: "There was a treasure. It was never discovered."
        • Example of No: "There was a treasure on a Greek island. It was never discovered."
    • Does it introduce new facts? (yes/no)
        Indicate:
          Yes if the SPRP hypothesis includes information that was not present in the source sentence;
          No if the SPRP hypothesis does not include information that was not present in the source sentence.
        Examples:
        • Source: "There was a treasure, which was never discovered."
        • Example of Yes: "There was a treasure on a British island. It was never discovered."
        • Example of No: "There was a treasure. It was never discovered."
    • Does it have splits in the wrong place or no split at all? (yes/no)
        Indicate:
          Yes if the SPRP hypothesis includes sentence splits in the wrong places or consists of only one sentence;
          No if the SPRP hypothesis does not include sentence splits in the wrong places and contains at least two separate sentences;
        Examples:
        • Source: "There was a treasure on a Greek island, which was never discovered and remained there forever."
        • Examples of Yes:
          "There was a treasure. On a Greek island it was never discovered. It remained there forever."
          "There was a treasure on a Greek island, which was never discovered and remained there forever."
        • Examples of No:
          "There was a treasure on a Greek island. It was never discovered. It remained there forever."
          "There was a treasure on a Greek island. It was never discovered and remained there forever."
    • Should some of its sentences be further split? (yes/no)
        Indicate:
          Yes if the SPRP hypothesis could be further split and rephrased into more sentences.
          No if the SPRP hypothesis could not be further split and rephrased into more sentences.
        Examples:
        • Source: "There was a treasure on a Greek island, which was never discovered and remained there forever."
        • Examples of Yes:
          "There was a treasure on a Greek island. It was never discovered and remained there forever."
          "There was a treasure on a Greek island, which was never discovered. It remained there forever."
        • Example of No: "There was a treasure on a Greek island. It was never discovered. It remained there forever."
  • NOTE: Some of the source sentences contain errors, of the types listed below. These cases should be evaluated as if the source sentence was a correct complex sentence: [^15] These cases had eluded our initial manual review of the data sampled from Wiki-BM. There were 10 such cases in total in the evaluation, and only 2 for the most problematic first case.
    • The source sentence actually contains two sentences, with missing space around the sentence separator, e.g., "xxx.They xxx".
    • There is a missing comma, e.g. "They are representatives from Tucson, Arizona who xxx".
    • There is a missing preposition, e.g. "The plan seemed work".
Appendix F Initial Multilingual Results
Our results in this work were based on English, for which different LLM variants and datasets were available. As a first approximation in other languages, in this section we provide results for German, French and Spanish, using a multilingual BLOOM 7.1B model [^16] https://huggingface.co/bigscience/bloom-7b1 Scao et al. (2022) to ensure better coverage than with the English-centric Pythia model. We also include English results with the BLOOM model, for comparison purposes.
Language Train Test BESC BLEU SARI add keep del CPL
English DeSSE DeSSE 89.05 82.00 69.56 45.51 92.42 70.74 94.05
English BiSECT BiSECT 76.18 41.50 42.10 14.90 64.45 46.95 81.72
French DeSSE DeSSE 89.30 65.00 53.52 24.79 81.18 54.59 83.93
French BiSECT BiSECT* 80.61 27.20 38.45 3.88 56.55 54.92 80.21
French BiSECT BiSECT 80.60 32.60 34.89 8.28 55.50 40.87 72.55
Spanish DeSSE DeSSE 90.67 67.10 54.30 23.95 83.28 55.67 87.70
Spanish BiSECT* BiSECT* 83.02 39.30 30.00 5.46 60.16 24.38 57.03
German DeSSE DeSSE 87.62 54.80 44.75 16.47 72.85 44.93 72.22
German BiSECT* BiSECT* 78.19 26.80 26.02 1.67 47.92 28.47 63.13
Table 10: Results in English, German, French and Spanish with the Bloom-7B1 model. Train and test datasets with missing diacritics are indicated with an asterisk. BertScore results are only normalised for English.
Since the DeSSE datasets were only available in English, we used an NLLB MT model [^17] https://huggingface.co/facebook/nllb-200-3.3B Costa-jussà et al. (2022) to translate the corpora into each language. Although this approximation can introduce translation errors in the process, it was the only available alternative for the DeSSE scenario.
The BiSECT corpora were available for all three of the aforementioned languages, but suffered from key issues: to the exception of the French training corpus, all datasets were missing diacritics, which are critical in these languages. To partially cope with this issue, for French we created an additional test set by randomly extracting 2,000 pairs from the original training dataset; for the other languages, we report results using the faulty datasets, as there was no alternative in these cases. Train and dev size were the same as for English, whereas the test sets contained 1036, 2018 and 735 pairs for French, Spanish and German, respectively.
The results for these experiments are summarised in Table 10. The English results obtained with the BLOOM model were similar to those obtained with the Pythia model (see Table 3), indicating that the overall SPRP ability does not seem impacted by variance between fine-tuned models of similar overall architecture and parameter size.
On the machine-translated DeSSE datasets, the results were high overall across metrics for Spanish, French, German, in descending accuracy order, although below those obtained in English, presumably as a result of the comparatively smaller volumes of LLM training data in languages other than English. On BiSECT, of note are the results obtained with models trained and tested on datasets that lacked diacritics. These models still generated output that included correct word forms with diacritics, the underlying model having been trained on correct data. The scores obtained against test sets without diacritics were thus negatively affected, and all models but the French one were trained on faulty data, an additional artificial constraint. The results in French on the test set with diacritics are below English but within range overall. They are also below the results on the faulty test set in terms of SARI and compliance, but these differences may be due to the non-faulty test set being twice the size of the original one.
Overall, these results tend to indicate strong potential of multilingual LLMs for the SPRP task, factoring the lack of, or specific issues with, existing datasets in languages other than English.